OpenStudy (anonymous):

give the slope and y-intercept of each line whose equation is given. then graph the linear function. Y=-3/5(x)+7

OpenStudy (anonymous):

slope=-3/5 y-intercept=7

OpenStudy (anonymous):

y=mx+b is slope intercept form of an equation m is the slope and b is the y-intercept

OpenStudy (anonymous):

to graph put a point at (0,7) --> 7 on the y-axis and then count down 3 and to the right 5 and put a point at (5,4), do that again and if you can put a point at (10,1) and draw a line through the points